# RUBIES: a complete census of the bright and red distant Universe with JWST/NIRSpec

### v1

@@
Although unified by having red colours over , the new sources discovered with the NIRCam instrument onboard JWST (Rieke et al., 2023) have highly heterogeneous morphologies (e.g. Nelson et al., 2023; Pérez-González et al., 2023; Labbe et al., 2023), suggestive of multiple classes of objects with different formation paths. At the highest redshifts, a red colour (with respect to HST) typically reflects the Lyman break in the rest-frame UV (e.g. Madau et al., 1996; Steidel et al., 1996, 2003) coupled with either strong emission lines or possibly a Balmer break at rest-frame optical wavelengths (e.g. Eyles et al., 2005; Labbé et al., 2010). Searches designed for these types of spectral energy distributions (SED) have rapidly yielded a vast number of candidate galaxies at , and beyond (e.g. Castellano et al., 2022; Finkelstein et al., 2023; Atek et al., 2022; Donnan et al., 2022; Adams et al., 2022; Naidu et al., 2022). Several of these candidates were significantly brighter than anticipated and suggested to be extremely massive galaxies, reaching stellar masses of before the Universe is 800 Myr old (Labbé et al., 2023). The abundance and masses of these systems have sparked a debate whether these findings are consistent with the standard cosmological model (Boylan-Kolchin, 2023).
@@
At redshifts , red sources that are not detected at but luminous at may be strongly obscured by dust. Mid- and far-infrared missions as well as ground-based sub-millimetre facilities had previously uncovered a population of sources that are extremely luminous at such long wavelengths, but often faint or undetected with HST, especially at higher redshifts (e.g. Franco et al., 2018; Wang et al., 2019; Casey et al., 2019; Williams et al., 2019; Manning et al., 2022): these sub-millimetre galaxies (SMGs) are typically at a redshift of , and are thought to be massive galaxies with extremely high star formation rates (for a review, see Casey et al., 2014; Hodge & da Cunha, 2020). Thanks to the improved sensitivity of JWST, we are now able to detect near-infrared emission from such SMGs out to (Herard-Demanche et al., 2023; Sun et al., 2024), and also extend this population to lower luminosities (Price et al., 2023). The newly discovered population of extremely red sources likely contributes significantly to the stellar mass budget of the high-redshift Universe (e.g. Nelson et al., 2023; Fudamoto et al., 2022; Barrufet et al., 2023; Gottumukkala et al., 2024; Pérez-González et al., 2023; Xiao et al., 2023; Weibel et al., 2024b; Williams et al., 2024).
@@
In contrast with this population of highly star-forming galaxies, a large number of photometric candidate massive quiescent galaxies have been identified out to (Carnall et al., 2023a; Long et al., 2024; Valentino et al., 2023; Pérez-González et al., 2023). The broad-band SEDs of these systems are consistent with strong Balmer breaks, resulting in a red colour (). Indeed, spectroscopic follow-up with JWST/NIRSpec has now confirmed the presence of old stellar populations in several of these systems at (e.g. Carnall et al., 2023b; de Graaff et al., 2024b), with the highest redshift massive quiescent galaxy discovered at (Weibel et al., 2024a). The existence of such objects is surprising: the formation of massive () galaxies at simultaneously requires rapid mass assembly in the first Gyr, and cessation of star formation in an epoch where the star formation activity in galaxies is typically only increasing. The great abundance of massive quiescent galaxies at these high redshifts would pose a challenge for many galaxy formation models (Valentino et al., 2023).
@@
Lastly, a mysterious sample of extremely compact red sources is ill-described by all of the above classes of objects. An apparently characteristic feature is a ‘v-shaped’ SED (e.g. Furtak et al., 2023; Barro et al., 2024; Labbe et al., 2023), i.e. a blue rest-frame UV continuum and red rest-frame optical continuum, which has proved challenging to model with many standard SED fitting codes (e.g. Killi et al., 2023; Wang et al., 2024a). With photometric redshifts ranging from to , the nature of these sources remains highly debated. Some are likely to be cool dwarf stars in the Milky Way (e.g. Burgasser et al., 2024; Hainline et al., 2024; Holwerda et al., 2024), while the first spectroscopic measurements for others have unveiled broad Balmer lines suggestive of accreting black holes (e.g. Kocevski et al., 2023; Harikane et al., 2023; Matthee et al., 2024; Greene et al., 2024). If the photometric redshifts are correct and the emission originates from active galactic nuclei (AGN), then these sources may reflect the early formation of massive black holes in high-redshift galaxies, challenging models of black hole growth (Greene et al., 2024). However, if the SED is instead dominated by stars, some of these objects may represent the most massive systems in the high-redshift Universe, forming the likely progenitors of early-type galaxies at (Labbé et al., 2023; Baggen et al., 2023; Akins et al., 2024; Wang et al., 2024b).
@@
The limiting factor in understanding the nature of these different bright and red sources in the early Universe is the coarse wavelength sampling from broadband photometry alone. Spectroscopy at near-infrared wavelengths is crucial to characterise the intrinsic shape of the SEDs and the presence of strong emission lines that can be degenerate with continuum breaks in broadband photometry. Multi-object spectroscopy with the NIRSpec instrument (Jakobsen et al., 2022; Böker et al., 2023) has proved extremely powerful thus far: even at modest depths, early spectroscopic programmes have confirmed the redshifts of over a dozen galaxies, revealed a great abundance of emission lines, as well as continuum emission in galaxies out to (e.g. Curti et al., 2023; Curtis-Lake et al., 2023; Roberts-Borsani et al., 2023; Fujimoto et al., 2023a, b; Wang et al., 2023).
@@
Large spectroscopic programmes in Cycle 1 have predominantly prioritised the search for the highest redshift galaxies, which typically are selected to have SEDs consistent with Lyman breaks with blue UV slopes (e.g. the NIRSpec guaranteed time observations (GTO) programmes and early release science programmes; Eisenstein et al. 2023; Maseda et al. 2024; Finkelstein et al. 2023; Treu et al. 2022). Moreover, many of these spectroscopic targets were selected from HST imaging catalogs. Because very red sources were not present in the photometric catalogs or not prioritised in the mask design procedure, the total number of such sources with follow-up spectroscopic observations has been extremely limited thus far.
@@ 2.1 Image data
In JWST Cycle 1, the EGS formed the focus of the Cosmic Evolution Early Release Science Survey (CEERS; PID 1345; PI: Finkelstein). The NIRCam imaging obtained in this programme spans 7 filters (F115W, F150W, F200W, F277W, F356W, F410M and F444W), reaching a point source depth (in a circular aperture of radius ) of 28.6 mag in the F444W filter over an area of approximately 80 arcmin2 (Bagley et al., 2023; Finkelstein et al., 2023). In addition, smaller sets of imaging data from programmes 2279 (PI: Naidu), 2514 (PIs: Williams & Oesch) and 2750 (PI: Arrabal Haro) add depth and area to some of the above filters. Finally, although not publicly available at the time of target selection, F090W imaging from programme 2234 (PI: Bañados; Khusanova et al. in prep.) covers the full CEERS footprint, which we use for flux calibration of the RUBIES spectra.
@@ 2.3 Parent catalogue
The parent catalogue of RUBIES was largely constructed from the source catalogues (version 7.2) that are publicly available on the DJA. These catalogues were created by performing source detection on an inverse variance weighted stack of the NIRCam F277W, F356W and F444W mosaics with SEP (Barbary, 2016), a Python implementation of SourceExtractor (Bertin & Arnouts, 1996), and subsequently measuring photometry for the detected sources in circular apertures of radius for all available bands. The aperture fluxes were rescaled to ‘total’ fluxes using the ratio between the Kron aperture flux and the circular aperture flux measured from the detection image. Importantly, these measurements do not account for variations in the point spread function (PSF) as a function of wavelength.
Photometric redshifts were estimated using eazy (Brammer et al., 2008) with the agn_blue_sfhz_13 template set and without any priors. This template set is optimised for a broad redshift range, and includes templates of emission line dominated sources, as well as an empirical template of a compact red AGN designed to roughly match the source from Killi et al. (2023). The fits were run using an iterative estimation of zero-point offsets for each filter: although the absolute flux calibration of NIRCam (Gordon et al., 2022) is currently better than for the filters used here333https://jwst-docs.stsci.edu/jwst-calibration-status/nircam-calibration-status/nircam-imaging-calibration-status, these empirical zero-point offsets effectively apply an average correction designed to somewhat compensate for the different PSFs of the images that were not convolved to a common PSF before the aperture photometry was performed.
@@ 2.4 Target prioritisation
Red sources. We selected all sources with and , where the colour was measured in a circular aperture of radius and a upper limit was used in case of non-detection in the F150W filter, and the upper limit on the magnitude was chosen to yield a marginal detection of continuum emission with the PRISM (pix-1). As we may expect template fitting to fail for very rare sources, we did not use any photometric redshift information for the selection (and we indeed find a high outlier fraction for these sources, as discussed in Section 3.2). Image cutouts of all sources in the DJA catalogue meeting these criteria were visually inspected, regardless of photometric quality flag, to check whether the sources are real or artefacts. This yielded 1269 sources across both fields.
Bright high-redshift sources. We selected sources with , probability and from either the DJA or the W24 catalogues (i.e. duplicate sources only need to meet these criteria in one catalogue to be selected). We further required that sources are covered by all available JWST broad filters, have a signal-to-noise ratio in the stacked long wavelength mosaic and are undetected () in filters below (F435W, F606W, and F814W or F090W where available). For the W24 catalogues we used photometry measured from smaller circular apertures of radius . The SEDs and image cutouts were visually inspected for all sources to assess whether the source is (i) real or an artefact and (ii) likely to be at high redshift. Sources were inspected by three reviewers (AdG, AW, PO) and selected as a good high-redshift candidate through a simple majority, resulting in a total of 868 objects.
@@ 2.5.3 Spectroscopic completeness
In total, RUBIES targets sources in both the PRISM and G395M observations. Approximately 300 of these targets are red, and 200 are bright high-redshift candidates (as defined in Section 2.4). An additional sources are observed with only the G395M disperser. Figure 3 shows the redshift, F444W and distributions of the selected targets in comparison to the parent sample, as well as the ratio between the target sample and parent sample. RUBIES is clearly strongly biased toward red sources, demonstrating the success of our mask design procedure. Despite what the survey name suggests, however, the majority of RUBIES targets are relatively blue: this reflects the fact that the vast majority of sources in the parent sample are blue, and these sources form the census sample that is critical to place the rare red sources into the context of the full galaxy population.
@@
Similarly, in Figure 5 we show the completeness as a function of colour and magnitude. As before, we distinguish between the PRISM and G395M masks, as well as the completeness computed for the RUBIES area and full NIRCam area. We further differentiate between the full survey and sources with . These figures demonstrate that RUBIES reaches very high completeness () for the reddest sources, especially for sources with , even though the photometric redshift was not explicitly included in the target prioritisation for the reddest sources (Section 2.4). As in Figure 4, comparing the two different measures of completeness shows that RUBIES is biased toward redder sources as the result of our pointing location optimisation.
@@ 3 Data processing
Moreover, we use the multiple observations of the RUBIES targets (i.e. both PRISM and G395M observations) to assess the relative flux and wavelength calibration of our dataset. A similar analysis was previously performed by the NIRSpec GTO team (Bunker et al., 2023; D’Eugenio et al., 2024) for data from the JWST Advanced Deep Extragalactic Survey (JADES Eisenstein et al., 2023) reduced with the NIRSpec GTO pipeline. The large number of RUBIES targets and our multi-disperser observing strategy now enables such a characterisation also for GO data and the public pipeline msaexp.
@@ 3.2 Spectroscopic redshifts
We used the template fitting implemented in msaexp to estimate spectroscopic redshifts for the reduced PRISM spectra (with global background subtraction) with the same template set that was used for the photometric redshifts. This omits the sources that were only observed with the G395M disperser, which we defer to a future data release paper. All PRISM spectra and template fits were visually inspected to (i) assess the redshift fit and (ii) check for major data quality issues. In this visual inspection process the best-fit redshift from msaexp can be manually updated to a different redshift by the inspector, although for the majority of sources the best-fit redshift matches the inspected redshift.

### v2

@@
Although unified by having red colours over , the new sources discovered with the NIRCam instrument onboard JWST (Rieke et al., 2023) have highly heterogeneous morphologies (e.g. Nelson et al., 2023; Pérez-González et al., 2023; Labbe et al., 2025), suggestive of multiple classes of objects with different formation paths. At the highest redshifts, a red colour (with respect to HST) typically reflects the Lyman break in the rest-frame UV (specifically, the spectral break at the Lyman limit of at , and the Lyman- break due to absorption by the neutral intergalactic medium at ; e.g. Madau et al. 1996; Steidel et al. 1996; Giavalisco 2002; Steidel et al. 2003) coupled with either strong emission lines or possibly a Balmer break at rest-frame optical wavelengths (e.g. Eyles et al., 2005; Labbé et al., 2010). Searches designed for these types of spectral energy distributions (SED) have rapidly yielded a vast number of candidate galaxies at , and beyond (e.g. Castellano et al., 2022; Finkelstein et al., 2023; Atek et al., 2022; Donnan et al., 2022; Adams et al., 2022; Naidu et al., 2022). Several of these candidates were significantly brighter than anticipated and suggested to be extremely massive galaxies, reaching stellar masses of before the Universe is 800 Myr old (Labbé et al., 2023). The abundance and masses of these systems have sparked a debate whether these findings are consistent with the standard cosmological model (Boylan-Kolchin, 2023) , and whether the redshift and mass estimates themselves are correct (e.g. Endsley et al., 2023; Kocevski et al., 2023).
@@
At redshifts , red sources that are not detected at but luminous at may be strongly obscured by dust. Mid- and far-infrared missions as well as ground-based sub-millimetre facilities had previously uncovered a population of sources that are extremely luminous at such long wavelengths, but often faint or undetected with HST, especially at higher redshifts (e.g. Franco et al., 2018; Wang et al., 2019; Casey et al., 2019; Williams et al., 2019; Manning et al., 2022): these sub-millimetre galaxies (SMGs) are typically at a redshift of , and are thought to be massive galaxies with extremely high star formation rates (for a review, see Casey et al., 2014; Hodge & da Cunha, 2020). Thanks to the improved sensitivity of JWST, we are now able to detect near-infrared emission from such SMGs out to (Herard-Demanche et al., 2025; Sun et al., 2024a), and also extend this population to lower luminosities (Price et al., 2025). The newly discovered population of extremely red sources likely contributes significantly to the stellar mass budget of the high-redshift Universe (e.g. Nelson et al., 2023; Fudamoto et al., 2022; Barrufet et al., 2023; Gottumukkala et al., 2024; Pérez-González et al., 2023; Xiao et al., 2024; Weibel et al., 2024b; Williams et al., 2024).
@@
In contrast with this population of highly star-forming galaxies, a large number of photometric candidate massive quiescent galaxies have been identified out to (Carnall et al., 2023a; Long et al., 2024; Valentino et al., 2023; Pérez-González et al., 2023). The broad-band SEDs of these systems are consistent with strong Balmer breaks, resulting in a red colour (). Indeed, spectroscopic follow-up with JWST/NIRSpec has now confirmed the presence of old stellar populations in several of these systems at (e.g. Carnall et al., 2023b; de Graaff et al., 2025), with the highest redshift massive quiescent galaxy discovered at (Weibel et al., 2024a). The existence of such objects is surprising: the formation of massive () galaxies at simultaneously requires rapid mass assembly in the first Gyr, and cessation of star formation in an epoch where the star formation activity in galaxies is typically only increasing. The great abundance of massive quiescent galaxies at these high redshifts would pose a challenge for many galaxy formation models (Valentino et al., 2023).
@@
Lastly, a mysterious sample of extremely compact red sources is ill-described by all of the above classes of objects. An apparently characteristic feature is a ‘v-shaped’ SED (e.g. Furtak et al., 2023; Barro et al., 2024; Labbe et al., 2025), i.e. a blue rest-frame UV continuum and red rest-frame optical continuum, which has proved challenging to model with many standard SED fitting codes (e.g. Killi et al., 2023; Wang et al., 2024a). With photometric redshifts ranging from to , the nature of these sources remains highly debated. Some are likely to be cool dwarf stars in the Milky Way (e.g. Burgasser et al., 2024; Hainline et al., 2024; Holwerda et al., 2024), while the first spectroscopic measurements for others have unveiled broad Balmer lines suggestive of accreting black holes (e.g. Kocevski et al., 2023; Harikane et al., 2023; Matthee et al., 2024; Greene et al., 2024). If the photometric redshifts are correct and the emission originates from active galactic nuclei (AGN), then these sources may reflect the early formation of massive black holes in high-redshift galaxies, challenging models of black hole growth (Greene et al., 2024). However, if the SED is instead dominated by stars, some of these objects may represent the most massive systems in the high-redshift Universe, forming the likely progenitors of early-type galaxies at (Labbé et al., 2023; Baggen et al., 2023; Akins et al., 2024; Wang et al., 2024b).
@@
The limiting factor in understanding the nature of these different bright and red sources in the early Universe is the coarse wavelength sampling from broadband photometry alone. Spectroscopy at near-infrared wavelengths is crucial to characterise the intrinsic shape of the SEDs and the presence of strong emission lines that can be degenerate with continuum breaks in broadband photometry. Multi-object spectroscopy with the NIRSpec instrument (Jakobsen et al., 2022; Böker et al., 2023) has proved extremely powerful thus far: even at modest depths, early spectroscopic programmes have confirmed the redshifts of over a dozen galaxies, revealed a great abundance of emission lines, as well as continuum emission in galaxies out to (e.g. Curti et al., 2023; Curtis-Lake et al., 2023; Roberts-Borsani et al., 2023; Fujimoto et al., 2023, 2024; Wang et al., 2023; Arrabal Haro et al., 2023a, b).
@@
Large spectroscopic programmes in Cycle 1 have predominantly prioritised the search for the highest redshift galaxies, which typically are selected to have SEDs consistent with Lyman breaks with blue UV slopes (e.g. the NIRSpec guaranteed time observations (GTO) programmes and early release science programmes; Eisenstein et al. 2023; Maseda et al. 2024; Finkelstein et al. 2023; Treu et al. 2022). Moreover, many of these spectroscopic targets were selected from HST imaging catalogues. Because very red sources were not present in the photometric catalogues or not prioritised in the mask design procedure, the total number of such sources with follow-up spectroscopic observations has been extremely limited thus far.
@@ 2.1 Image data
In JWST Cycle 1, the EGS formed the focus of the Cosmic Evolution Early Release Science Survey (CEERS; PID 1345, PI: Finkelstein; Finkelstein et al. 2025). The NIRCam imaging obtained in this programme spans 7 filters (F115W, F150W, F200W, F277W, F356W, F410M and F444W), reaching a point source depth (in a circular aperture of radius ) of 28.6 mag in the F444W filter over an area of approximately 80 arcmin2 (Bagley et al., 2023; Finkelstein et al., 2023). In addition, smaller sets of imaging data from programmes 2279 (PI: Naidu), 2514 (PIs: Williams & Oesch; Williams et al. 2025) and 2750 (PI: Arrabal Haro) add depth and area to some of the above filters. Finally, although not publicly available at the time of target selection, F090W imaging from programme 2234 (PI: Bañados; Khusanova et al. in prep.) covers the full CEERS footprint, which we use for flux calibration of the RUBIES spectra.
@@ 2.3 Parent catalogue
The parent catalogue of RUBIES was largely constructed from the source catalogues (version 7.2) that are publicly available on the DJA. These catalogues were created by performing source detection on an inverse variance weighted stack of the NIRCam F277W, F356W and F444W mosaics with SEP (Barbary, 2016), a Python implementation of SourceExtractor (Bertin & Arnouts, 1996), and subsequently measuring photometry for the detected sources in circular apertures of radius for all available bands. Uncertainties on these aperture fluxes were measured from the weight images, by summing the pixel variances in the same apertures. We note that we chose an aperture of radius , as this matches the effective radii of galaxies at the median photometric redshift of our survey (; e.g. Kartaltepe et al. 2023; Ormerod et al. 2024; Sun et al. 2024b). The aperture fluxes were rescaled to ‘total’ fluxes using the ratio between the Kron aperture flux and the circular aperture flux measured from the detection image. Importantly, these measurements do not account for variations in the point spread function (PSF) as a function of wavelength.
Photometric redshifts were estimated using eazy (Brammer et al., 2008) with the agn_blue_sfhz_13 template set and without any priors. This template set is optimised for a broad redshift range, and includes templates of emission line dominated sources, as well as an empirical template of a compact red AGN designed to roughly match the source from Killi et al. (2023). The redshift fits were run using an iterative estimation of zero-point offsets for each filter, as described in Whitaker et al. (2011) and Skelton et al. (2014). Briefly, this algorithm computes the residual between the observed and best-fit model photometry for each filter (keeping F277W as a reference point); the average zero point offset (per filter) is then computed from all objects in the catalogue. The redshift fitting is repeated with these new zero point estimates, in order to iteratively minimise the residuals for all filters. Although the absolute flux calibration of NIRCam (Gordon et al., 2022) is currently better than for the filters used here333https://jwst-docs.stsci.edu/jwst-calibration-status/nircam-calibration-status/nircam-imaging-calibration-status, we find that the DJA aperture photometry and PSF-matched aperture photometry of W24 differ by an approximately constant offset (of up to mag for short wavelengths), with secondary scatter due to source morphology and colour gradients. These empirical zero-point offsets therefore effectively apply an average correction that (partially) compensates for the different PSFs of the images that were not convolved to a common PSF before the aperture photometry was performed.
@@ 2.4 Target prioritisation
Red sources. We selected all sources with and , where the colour was measured in a circular aperture of radius and a upper limit was used in case of non-detection in the F150W filter. The upper limit on the magnitude () was chosen to yield a marginal detection of continuum emission within 48 min of PRISM observations (pix-1), which we estimated based on extensive simulations with the JWST Exposure Time Calculator using realistic galaxy size distributions. As we may expect template fitting to fail for very rare sources, we did not use any photometric redshift information for the selection (and we indeed find a high outlier fraction for these sources, as discussed in Section 3.2). Image cutouts of all sources in the DJA catalogue meeting these criteria were visually inspected, regardless of photometric quality flag, to check whether the sources are real or artefacts. This yielded 1269 sources across both fields.
Bright high-redshift sources. We selected sources with , probability and from either the DJA or the W24 catalogues (i.e. duplicate sources only need to meet these criteria in one catalogue to be selected). We further required that sources are covered by all available JWST broad filters, have a signal-to-noise ratio in the stacked long wavelength mosaic and are undetected () in filters below (F435W, F606W, and F814W or F090W where available). For the W24 catalogues we used (PSF-matched) photometry measured from smaller circular apertures of radius . The SEDs and image cutouts were visually inspected for all sources to assess whether the source is (i) real or an artefact and (ii) likely to be at high redshift. Sources were inspected by three reviewers (AdG, AW, PO) and selected as a good high-redshift candidate through a simple majority, resulting in a total of 868 objects.
@@ 2.5.3 Spectroscopic completeness
In total, RUBIES targets 2901 sources in both the PRISM and G395M observations. Approximately 300 of these targets are red, and 200 are bright high-redshift candidates (as defined in Section 2.4). An additional sources are observed with only the G395M disperser. Figure 3 shows the redshift, F444W and distributions of the selected targets in comparison to the parent sample, as well as the ratio between the target sample and parent sample. RUBIES is clearly strongly biased toward red sources, demonstrating the success of our mask design procedure. Despite what the survey name suggests, however, the majority of RUBIES targets are relatively blue: this reflects the fact that the vast majority of sources in the parent sample are blue, and these sources form the census sample that is critical to place the rare red sources into the context of the full galaxy population.
@@
Similarly, in Figure 5 we show the completeness as a function of colour and magnitude. As before, we distinguish between the PRISM and G395M masks, as well as the completeness computed for the RUBIES area and full NIRCam area. We further differentiate between the full survey and sources with . These figures demonstrate that RUBIES reaches very high completeness () for the reddest sources, especially for sources with , even though the photometric redshift was not explicitly included in the target prioritisation for the reddest sources (Section 2.4). As in Figure 4, comparing the two different measures of completeness shows that RUBIES is biased toward redder sources as the result of our pointing location optimisation (which maximises the number of red sources observed across the survey, see Section 2.5.1).
@@ 3 Data processing
Moreover, we use the multiple observations of the RUBIES targets (i.e. both PRISM and G395M observations) to assess the relative flux and wavelength calibration of our dataset. A similar analysis was previously performed by the NIRSpec GTO team (Bunker et al., 2024; D’Eugenio et al., 2025) for data from the JWST Advanced Deep Extragalactic Survey (JADES Eisenstein et al., 2023) reduced with the NIRSpec GTO pipeline. The large number of RUBIES targets and our multi-disperser observing strategy now enables such a characterisation also for GO data and the public pipeline msaexp.
@@ 3.2 Spectroscopic redshifts
We used the least squares template fitting method implemented in msaexp to estimate spectroscopic redshifts for the reduced PRISM spectra (with global background subtraction) with the same template set that was used for the photometric redshifts. This omits the sources that were only observed with the G395M disperser, which we defer to a future data release paper. All PRISM spectra and template fits were visually inspected to (i) assess the redshift fit and (ii) check for major data quality issues. In this visual inspection process the best-fit redshift from msaexp can be manually updated to a different redshift by the inspector, although for the majority of sources the best-fit redshift matches the inspected redshift.